John is 8 years older than paul eight years to come their ages will be in the ratio of 5:4 respectively,find the age of each one of them
Answer by ikleyn(52799)

.
From the problem we know that in 8 years to come John's age will be 5x, while Paul age will be 4x, 

where x is the common measure of their ages at THAT time.


We are given also that the difference of their ages is (and, hence, always will be, until they both are alive) 8 years.


Hence,  5x - 4x = 8,   or  x = 8.


It means that in 8 years to come John's age will be 5*8 = 40 years, while Paul's age will be 4*8 = 32 years.


It implies that TODAY  John is 40-8 = 32 years old; Paul is 32-8 = 24 years old.    ANSWER

Solved.